OpenStudy (anonymous):

p(10)≈ 100e ^0.01(10)

OpenStudy (anonymous):

70

OpenStudy (anonymous):

where on earth did you get 70 from?

OpenStudy (anonymous):

0.01*10 = 0.1 e^0.1 is approximately 1.105 100*1.105 = 110.5

OpenStudy (anonymous):

sorry but what does .01e^.01 mean ?

OpenStudy (anonymous):

ok so from 100*e^10(0.01). From bedmas we have to do the brackets first and that is 10*0.01 = 0.1 then now we got 100*e^0.1 now we have to do the exponents e^0.1 is approximately =1.105 so now we got 100*1.105 and now we must do multipication and that is approxiamately 110.5

OpenStudy (anonymous):

oh ok, but we use the 100 to multiply twice ? it was use to multiply 0.01 already, then we multiply again with the 1.105?

OpenStudy (anonymous):

log4